The length of a rectangle is 5m more than its width. If its perimeter is 60m, what are the dimensions of the rectangle
Found 2 solutions by Cromlix, ikleyn:
Answer by Cromlix(4381)

Hi there,
Make width = 'x'
Length = x + 5
Perimeter of a rectangle
= 2(width) + 2(length)
60m = 2(x) + 2(x + 5)
60m = 2x + 2x + 10
Collect like terms:
2x + 2x = 60 - 10
4x = 50
x = 12.5
Width = 12.5 metres
Length = 17.5 metres
Hope this helps. :-)

Answer by ikleyn(52781)